OpenStudy (anonymous):

Use the function below to find F(4) F(x)=5•(1/3)^x A.20/3 B.20/81 C.5/81 D.5/12

OpenStudy (danjs):

\[F() = 5(\frac{ 1 }{ 3 })^4\] simpolify

OpenStudy (danjs):

F(4)

OpenStudy (danjs):

\[[\frac{ a }{ b }]^c = \frac{ a^c }{ b^c }\]

OpenStudy (danjs):

parenthesis/power part fitst, then multiply it by 5,, order of operations

OpenStudy (danjs):

\[5*\frac{ 1^4 }{ 3^4 }\]

OpenStudy (anonymous):

Thanks!